OpenStudy (mejia941):

Select steps that could be used to solve the equation 1 + 3x = -x + 4.

OpenStudy (brainzonly):

Combine like terms (I might be slow, i'm working in two qs at once)

OpenStudy (brainzonly):

which means following the Golden Rule of Algebra: subtract 1 from both sides of the equation: 1-1+3x= -x+4-1 3x=-x+3 Add x to both sides: 3x+x=-x+x+3 4x=3 x=3/4

OpenStudy (brainzonly):

have a good day